# A stochastic preconditioned Douglas-Rachford splitting method for saddle-point problems

### v2

@@ Abstract
In this article, we propose and study a stochastic preconditioned Douglas–Rachford splitting method to solve saddle-point problems that have separable dual variables. We prove the almost sure convergence of the iteration sequences in Hilbert spaces for a class of convex-concave and nonsmooth saddle-point problems. We also provide the sublinear convergence rate for the ergodic sequence concerning the expectation of the restricted primal-dual gap functions. Numerical experiments show the high efficiency of the proposed stochastic preconditioned Douglas–Rachford splitting methods.
@@
Douglas-Rachford splitting, stochastic optimization, saddle-point problems, almost sure convergence, preconditioning techniques.
MSC codes.
@@ 1 Introduction
The Douglas–Rachford (henceforth abbreviated as DR) splitting method is an iterative splitting method for finding a zero point of the sum of maximal set-valued monotone operators, which is a powerful optimization tool for convex optimization and has plenty of applications [3, 8, 14]. The method originated from splitting linear operators for solving heat conduction equations [13] and was extended to a more general nonlinear maximal monotone operator splitting framework [18, 23]. The convergence of the DR splitting method in Hilbert spaces can be found in [14] through a proximal point framework. The DR splitting method is essentially an implicit and unconditionally stable splitting algorithm for arbitrary step sizes. Solving implicit equations including linear equations can be computationally demanding for DR methods. In [8], within the generalized proximal-point framework, an efficient preconditioned DR framework with a convergence guarantee aiming at a specific class of convex-concave saddle-point problems is developed. It can incorporate the classical preconditioning techniques for linear systems into the nonlinear DR iterations. We refer to [7] for applications of preconditioned DR and [6] for its accelerated variants.
Due to a wide range of applications in various fields including big data science and machine learning, there are lots of studies on stochastic first-order algorithms [4, 11] and especially for saddle-point problems [1, 10, 11]. In this paper, we study a stochastic preconditioned DR splitting framework which includes the stochastic DR without preconditioners [11] as a special case for solving the following saddle-point problems with separable dual variables [1, 11].
@@
Compared to related work, our contributions are as follows. First, different from the stochastic first-order primal-dual method [10] or stochastic classical DR splitting method without preconditioners [11], preconditioners are introduced and the challenges in the convergence analysis are addressed. Actually, preconditioners already introduced challenges for the deterministic preconditioned DR and extra effort is necessary for the corresponding convergence analysis [8, 6]. For the stochastic preconditioned DR (henceforth abbreviated as SPDR), the almost sure convergence framework of the stochastic fixed point iteration based on firmly nonexpansive operators developed in [11] cannot be applied directly here. We circumvent these difficulties by analyzing the relation between the stochastic updates and the one-step deterministic updates based on the current stochastic iteration. We mainly employed the celebrated stochastic quasi-Fejér monotonicity [11, 15] and Opial’s lemma [2, 20]. Second, we show the SPDR splitting method has a $\mathcal{O}(1/k)$ rate for the ergodic sequence respecting restricted primal-dual gap functions where $k$ is the iteration number. Since taking the expectation of the supremum of the primal-dual gap is a very subtle issue [1], a detailed analysis is devoted to this part. Finally, we show the high efficiency of the proposed SPDR compared to state-of-art methods including the stochastic primal-dual method [10] with some numerical experiments.
The rest of the paper is organized as follows. In section 2, we give a brief introduction of the preconditioned deterministic DR and the proposed stochastic preconditioned DR splitting method along with some necessary analysis tools. In section 3, we discuss the almost sure convergence of the proposed SPDR splitting method to a fixed point using stochastic quasi-Fejér monotonicity [11] and Opial’s lemma [2]. In section 4, we show the sublinear convergence rate of ergodic sequences respecting the expectation of the restricted primal-dual gap. In section 5, we extend these results to the stochastic and preconditioned DR method for quadratic-linear problems (henceforth abbreviated as SPDRQ). In section 6, we present detailed numerical experiments for total generalized variation (TGV) regularized problems to show the competitive performance of the proposed stochastic and preconditioned DR methods. Finally, we give some conclusions in section 7.
@@ 2 Preconditioned DR framework
where $\sigma$ and $\tau$ are arbitrary positive primal and dual step sizes, $N_{1}\geq I$, $N_{2}\geq I$ with $N_{1}:\mathbb{X}\rightarrow\mathbb{X}$ and $N_{2}:\mathbb{Y}\rightarrow\mathbb{Y}$ being bounded, linear and self-adjoint operators, and $x_{\operatorname{test}}^{k+1}$ and $y_{\operatorname{test}}^{k+1}$ are transitional variables. Denoting by $z=(x,y)\in\mathbb{X}\times\mathbb{Y}$ and $\bar{z}=(\bar{x},\bar{y})\in\mathbb{X}\times\mathbb{Y}$, the iteration (PDR) with $N_{1}=I$, $N_{2}=I$ and $\sigma=\tau$ becomes the original DR splitting method [14, 18] based on the following maximal monotone operator splitting by the first-order optimality conditions of (1.2),
$0\in Az+Bz,\ \ \text{with}\quad A=\begin{pmatrix}0&\mathcal{K}^{*}\\ -\mathcal{K}&0\end{pmatrix},\quad B=\begin{pmatrix}\partial\mathcal{F}&0\\ 0&\partial\mathcal{G}\end{pmatrix}.$
Note that the first two equations of (PDR) are implicit. However, if we choose $N_{2}=I$ and an operator $M$ such that
@@
where $M:X\rightarrow X$ is linear, continuous, self-adjoint and satisfies $M\geq T:=I+\sigma\tau\mathcal{K}^{*}\mathcal{K}$, we can update $x^{k+1}$ first with the following classical preconditioned iteration
@@
and then obtain the linear update $y^{k+1}=\bar{y}^{k}+\tau\mathcal{K}x^{k+1}$ according to the second equation of (PDR) followed by $x^{k+1}$ [8, 7, 6]. Classical preconditioning techniques can be employed with (2.2) which can be seen as an iterative method for solving $Tx^{k+1}=b^{k}$. This linear equation arises from the linear update step in the original Douglas–Rachford splitting method where $N_{1}=I$ and $N_{2}=I$. Preconditioned DR splitting can bring high efficiency with great flexibility together with convergence guarantees using finitely many iterations of any feasible preconditioner [8, 7, 6]. A preconditioner is called feasible if $M\geq T$ as in (2.2) includes a lot of classical efficient preconditioners [8]. Let us denote by $u:=(x,y,\bar{x},\bar{y})\in\mathbb{U}:=(\mathbb{X}\times\mathbb{Y})^{2}$ and introduce an associated positive semidefinite and possibly degenerate inner product as
@@
with $\mathbf{M}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\sigma}I,\tfrac{1}{\tau}I]$. Denoting by $\mathcal{T}:\mathbb{U}\rightarrow\mathbb{U}$ as the iteration operator that maps $u^{k}=(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$ to $u^{k+1}=(x^{k+1},y^{k+1},\bar{x}^{k+1},\bar{y}^{k+1})$ according to (PDR), we can rewrite the iteration (PDR) as
@@
Now, let us give several lemmas which are convenient for analyzing the stochastic PDR method in the sequel, whose proofs can be found in [8, 7, 6] and are omitted here.
Lemma 1.
@@
The iteration (PDR) for (1.2) possesses the following properties:
@@
    A point $u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})$ is a fixed point of (2.4) if and only if $(x^{*},y^{*})$ is a saddle point for (1.2) and $x^{*}+\sigma\mathcal{K}^{*}y^{*}=\bar{x}^{*},y^{*}-\tau\mathcal{K}x^{*}=\bar{y}^{*}$.
@@
$\mathcal{F}(x)=\langle\frac{1}{2}Qx-f,x\rangle,$
where $Q$ is a self-adjoint, linear, and bounded positive semi-definite operator and $f\in X$, we can get a more compact preconditioned Douglas–Rachford splitting method for quadratic problems, which is called PDRQ [8, 7, 6]. Denoting by $u=(x,y,\bar{y})\in\mathbb{U}_{Q}:=\mathbb{X}\times\mathbb{Y}\times\mathbb{Y}$, $N_{1}\geq 0$, $N_{2}\geq I$, the PDRQ iteration reads as,
@@
Moreover, if we choose $N_{2}=I$ and $N_{1}$ such that
@@
where $M_{Q}:X\rightarrow X$ is linear, continuous, self-adjoint and feasible meaning $M_{Q}\geq T_{Q}:=\sigma Q+\sigma\tau\mathcal{K}^{*}\mathcal{K}$, then we can update $x^{k+1}$ first with the following classical preconditioned iteration
@@
The linear update of $y^{k+1}=\bar{y}^{k}+\tau\mathcal{K}x^{k+1}$ is then according to the second line in (PDRQ) [8, 7, 6]. Denoting by
@@
Denoting by $\mathcal{T}_{Q}$ as the iteration operator associated with (PDRQ), we can rewrite the iteration (PDRQ) as
@@
Similarly, we have the following proposition and lemma for PDRQ.
@@
$\mathcal{L}(x_{\operatorname{test}}^{k+1},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k+1})$ $\leq\langle u^{k}-u^{k+1},u^{k+1}-u\rangle_{\mathbf{M}_{Q}}$
$=\frac{1}{2}(\|u^{k}-u\|_{\mathbf{M}_{Q}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}_{Q}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}_{Q}}^{2}).$ (2.11)
Similar to Lemma 2, we give the following lemma with omitted proof, which will also turn out useful for the almost sure convergence of stochastic (2) method introduced in the paper.
@@
The iteration (PDRQ) for (1.2) possesses the following properties:
@@
With these preparations for deterministic PDR (or PDRQ), let us now turn to our focus, to the stochastic PDR method.
2.1 The stochastic preconditioned DR (SPDR) method
Let us begin with the stochastic setting. Denote by $(\Omega,\mathfrak{F},\mathcal{P})$ the underlying probability space. Let $\{i_{0},i_{1},i_{2},\ldots\}$ be index selection variables, i.e., independent random variables on the latter space with values in $\{1,\ldots,n\}$ which are identically distributed. We denote the probability of selecting an index $i\in\{1,\ldots,n\}$ by $p_{i}$ where we assume $10<p_{i}<$ 1 for each $i$ and recall that $\sum_{i=1}^{n}p_{i}=1$. Introduce $P=\operatorname{Diag}[p_{1},p_{2},\ldots,p_{n}]$ and observe that $P^{-1}=\operatorname{Diag}[\frac{1}{p_{1}},\frac{1}{p_{2}},\ldots,\frac{1}{p_{n}}]$. A Borel random vector $y$ in a separable Hilbert space $\mathbb{H}$ is a measurable map from $\Omega$ into $\mathbb{H}$ where $\Omega$ is equipped with $\mathfrak{F}$ and $\mathbb{H}$ is equipped with the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by the open sets of $\mathbb{H}$ (see [11] or Chapter 2 of [17]). Further, $(\mathfrak{F}_{k})_{k\in\mathbb{N}}$ denotes the filtration generated by the indices $\{i_{0},i_{1},\ldots,i_{k}\}$, i.e., the $\sigma$-algebras generated by $i_{j}^{-1}\{i\}$ for $j=0,\ldots,k$ and $i=1,\ldots,n$ [22] (Chapter 7). Then, $\mathfrak{F}_{k}\subset\mathfrak{F}_{k+1}$ for each $k\in\mathbb{N}$. Henceforth, we denote by $\mathbb{E}^{k}:=\mathbb{E}[\cdot\mid\mathfrak{F}_{k}]$ the conditional expectation with respect to $\mathfrak{F}_{k}$, and by $\mathbb{E}$ the full expectation. Recall that if a Borel random vector $y$ is $\mathfrak{F}_{k}$-measurable, i.e., the preimages of Borel sets are in $\mathfrak{F}_{k}$, then $\mathbb{E}[y\mid\mathfrak{F}_{k}]=y$.
Let us first extend the preconditioned Douglas–Rachford splitting method to a stochastic setting. For arbitrary positive step sizes $\sigma$, $\tau$ and $(x^{0},y^{0},\bar{x}^{0},\bar{y}^{0})\in(\mathbb{B}_{1}\times\mathbb{B}_{2})^{2}$, $y^{0}_{\operatorname{test}}\in\mathbb{B}_{2}$, the stochastic PDR (SPDR) reads as follows.
$\left\{\begin{array}[]{ll}x^{k+1}=\left(N_{1}+\sigma\tau\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(\left(N_{1}-I\right)x^{k}+\bar{x}^{k}-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right),\\ y^{k+1}=\left\{\begin{array}[]{ll}N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}x^{k+1}]&i=i_{k},\\ y^{k}_{i}&\forall i\neq i_{k},\\ \end{array}\right.\\ x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}(I+\tau\partial\mathcal{G}_{i})^{-1}[2y^{k+1}_{i}-\bar{y}^{k}_{i}]&i=i_{k},\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k},\\ \end{array}\right.\\ \bar{x}^{k+1}=\bar{x}^{k}+(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}]-x^{k+1},\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\bar{y}_{i}^{k}+y_{\operatorname{test},i}^{k+1}-y_{i}^{k+1}&i=i_{k},\\ \bar{y}^{k}_{i}&\forall i\neq i_{k},\end{array}\right.\end{array}\right.$ (SPDR)
where $N_{1}\geq I$ and $N_{2}=\operatorname{Diag}[N_{2,1},\ldots,N_{2,n}]$ with each $N_{2,i}:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$ being self-adjoint, bounded and linear such that $N_{2,i}\geq I$. Although the update of $x^{k+1}$ in the above (SPDR) seems complicated, however, it is convenient for analysis and the preconditioned iteration (2.2) can be recovered from (SPDR) with $N_{2}=I$ and $N_{1}$ as in (2.1).
Next, suppose that the bounded set $\mathbb{B}:=\mathbb{B}_{1}\times\mathbb{B}_{2}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$ contains a saddle point. We give some distance functions including the following restricted primal-dual gap $\mathfrak{G}_{\mathbb{B}_{1}\times\mathbb{B}_{2}}$ associated with $\mathbb{B}$:

### v3

@@ Abstract
In this article, we propose and study a stochastic and relaxed preconditioned Douglas–Rachford splitting method to solve saddle-point problems that have separable dual variables. We prove the almost sure convergence of the iteration sequences in Hilbert spaces for a class of convex-concave and nonsmooth saddle-point problems. We also provide the sublinear convergence rate for the ergodic sequence concerning the expectation of the restricted primal-dual gap functions. Numerical experiments show the high efficiency of the proposed stochastic and relaxed preconditioned Douglas–Rachford splitting methods.
@@
Douglas–Rachford splitting, stochastic optimization, saddle-point problems, over-relaxation, almost sure convergence, preconditioning techniques.
MSCodes.
@@ 1 Introduction
The Douglas–Rachford (henceforth abbreviated as DR) splitting method is an iterative splitting method for finding a zero point of the sum of maximal set-valued monotone operators, which is a powerful optimization tool for convex optimization and has plenty of applications [4, 10, 17, 25, 19]. The method originated from splitting linear operators for solving heat conduction equations [18] and was extended to a more general nonlinear maximal monotone operator splitting framework [32, 41]. The convergence of the DR splitting method in Hilbert spaces can be found in [19] through a proximal point framework. We further refer to [11, 26, 27, 22] for general DR splitting methods for convex problems and [2, 33, 42, 31] for nonconvex problems. The DR splitting method is essentially an implicit and unconditionally stable splitting algorithm, which means the DR splitting method is convergent with arbitrary positive step sizes. Solving implicit equations including linear equations can be computationally demanding for DR methods. In [10], within the generalized proximal-point framework, an efficient preconditioned DR framework with a convergence guarantee aiming at a specific class of convex-concave saddle-point problems is developed. It can incorporate the classical preconditioning techniques for linear systems into the nonlinear DR iterations. We refer to [9] for applications of preconditioned DR and [7] for its accelerated variants.
Due to a wide range of applications in various fields including big data science and machine learning, there are a lot of studies on stochastic first-order algorithms [5, 15] and especially for saddle-point problems [1, 13, 15]. In this paper, we study a stochastic preconditioned DR splitting framework which includes the stochastic DR without preconditioners [15] as a special case for solving the following saddle-point problems with separable dual variables [1, 15]. We refer to [24] for the randomized DR methods aiming at solving linear systems.
@@
Compared to related work, our contributions are as follows. First, different from the stochastic first-order primal-dual method [13] or stochastic classical DR splitting method without preconditioners [15], preconditioners are introduced. The challenges in the convergence analysis are addressed for the proposed stochastic DR splitting method. Preconditioners already introduced challenges for the deterministic preconditioned DR and extra effort is necessary for the corresponding convergence analysis [10, 7]. Second, we incorporate a general over-relaxed strategy into the stochastic and preconditioned DR splitting method with different relaxation parameters on different variables. It is well known that over-relaxation techniques can accelerate the deterministic DR splitting algorithm [19, 17, 33, 8]. For the over-relaxed, stochastic, and preconditioned DR (hereafter abbreviated as SRPDR), the almost sure convergence framework of the stochastic fixed point iteration based on firmly nonexpansive operators developed in [15] cannot be applied directly here. We circumvent these difficulties by analyzing the relation between the stochastic updates and the one-step deterministic updates based on the current stochastic iteration. We mainly employed the celebrated stochastic quasi-Fejér monotonicity [15, 16, 20] and Opial’s lemma [35]. Third, we show the SRPDR splitting method has a $\mathcal{O}(1/K)$ rate for the ergodic sequence respecting restricted primal-dual gap functions where $K$ is the iteration number. Since taking the expectation of the supremum of the primal-dual gap concerning transitional variables is a very subtle issue [1], a detailed analysis is devoted to this part. Furthermore, we also show the SRPDR splitting method has a $\mathcal{O}(1/K)$ rate for primal error when Lipschitz continuity is satisfied for dual functions. Finally, we show the high efficiency of the proposed SRPDR compared to state-of-art methods including the stochastic primal-dual method [13] with some numerical experiments on synthetic and real datasets.
The rest of the paper is organized as follows. In section 2, we give a brief introduction of the preconditioned deterministic DR and the proposed stochastic, over-relaxed, and preconditioned DR splitting method along with some necessary analysis tools. In section 3, we discuss the almost sure convergence of the proposed SRPDR splitting method to a fixed point using stochastic quasi-Fejér monotonicity [15] and Opial’s lemma [35]. In section 4, we show the sublinear convergence rate of ergodic sequences respecting the expectation of the restricted primal-dual gap and the primal error. In section 5, we extend these results to the stochastic relaxed and preconditioned DR method for quadratic-linear problems (henceforth abbreviated as SRPDRQ). In section 6, we present detailed numerical experiments for total generalized variation (TGV) regularized problems and binary classification with real datasets to show the competitive performance of the proposed stochastic and preconditioned DR methods. Finally, we give some conclusions in section 7.
@@ 2 Preconditioned DR framework
where $\sigma$ and $\tau$ are arbitrary positive primal and dual step sizes, $N_{1}\succeq I$, $N_{2}\succeq I$ with $N_{1}:\mathbb{X}\rightarrow\mathbb{X}$ and $N_{2}:\mathbb{Y}\rightarrow\mathbb{Y}$ being bounded, linear and self-adjoint operators, and $x_{\operatorname{test}}^{k+1}$ and $y_{\operatorname{test}}^{k+1}$ are transitional variables. Denoting by $z=(x,y)\in\mathbb{X}\times\mathbb{Y}$ and $\bar{z}=(\bar{x},\bar{y})\in\mathbb{X}\times\mathbb{Y}$, the iteration PDR with $N_{1}=I$, $N_{2}=I$ and $\sigma=\tau$ becomes the original DR splitting method [19, 32]. It was based on the following maximal monotone operator splitting
$0\in Az+Bz,\ \ \text{with}\quad A=\begin{pmatrix}0&\mathcal{K}^{*}\\ -\mathcal{K}&0\end{pmatrix},\quad B=\begin{pmatrix}\partial\mathcal{F}&0\\ 0&\partial\mathcal{G}\end{pmatrix},$
by the first-order optimality conditions of (1.2). Note that the first two equations of PDR are implicit. However, if we choose $N_{2}=I$ and an operator $M$ such that
@@
where $M:X\rightarrow X$ is linear, continuous, self-adjoint and satisfies $M\succeq T:=I+\sigma\tau\mathcal{K}^{*}\mathcal{K}$, we can update $x^{k+1}$ first with the following classical preconditioned iteration
@@
and then obtain the linear update $y^{k+1}=\bar{y}^{k}+\tau\mathcal{K}x^{k+1}$ according to the second equation of PDR followed by $x^{k+1}$ [10, 9, 7]. Classical preconditioning techniques can be employed with (2.2) which can be seen as an iterative method for solving $Tx^{k+1}=b^{k}$. This linear equation arises from the linear update step in the original Douglas–Rachford splitting method where $N_{1}=I$ and $N_{2}=I$. Preconditioned DR splitting can bring high efficiency with great flexibility and convergence guarantees using finitely many iterations of any feasible preconditioner [10, 9, 7]. A preconditioner is called feasible if $M\succeq T$ as in (2.2) includes a lot of classical efficient preconditioners [10]. Let us denote by $u:=(x,y,\bar{x},\bar{y})\in\mathbb{U}:=(\mathbb{X}\times\mathbb{Y})^{2}$ and introduce an associated positive semidefinite and possibly degenerate inner product as
@@
with $\mathbf{M}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\sigma}I,\tfrac{1}{\tau}I]$. Denoting by $\mathcal{T}:\mathbb{U}\rightarrow\mathbb{U}$ as the iteration operator that maps $u^{k}=(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$ to $u^{k+1}=(x^{k+1},y^{k+1},\bar{x}^{k+1},\bar{y}^{k+1})$ according to PDR, we can rewrite the iteration PDR as
@@
Now, we present the following lemma and proposition, which are useful for analyzing the stochastic PDR method in the subsequent sections. The proofs for the case where $\sigma=\tau$ can be found in [10, 9], while the proofs for the case where $\sigma\neq\tau$ can be obtained easily.
Lemma 1 ([9, Lemma 2.1]).
@@
The equality in (2.6) comes from the polarization identity with positive semi-definite weight.
Next, let us introduce the relaxed preconditioned Douglas-Rachford splitting method,
$u^{k+1}=(I-I_{\varrho_{k}})u^{k}+I_{\varrho_{k}}\mathcal{T}u^{k},$ (2.7)
where $I_{\varrho_{k}}:=\textrm{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,x}I,\rho_{k,y}I]$. Throughout the paper, we assume the dimension of each identity operator $I$ depends on the corresponding variable. Suppose $\rho_{k}$, $\rho_{k,x}$, and $\rho_{k,y}$ are non-decreasing numbers with lower and upper limits $\rho_{l}$ and $\rho_{u}$, respectively, such that $20<\rho_{l}<\rho_{u}<$ 2. With $\mathcal{T}$ being the iteration operator of PDR given by (2.4), we then define $u^{k+1}_{t}=\mathcal{T}u^{k}:=(x^{k+1}_{t},y^{k+1}_{t},\bar{x}^{k+1}_{t},\bar{y}^{k+1}_{t})$. In the computation, we also introduce the following new transitional variables:
$\left\{\begin{array}[]{ll}x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2x_{t}^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=(I+\tau\partial\mathcal{G})^{-1}[2y_{t}^{k+1}-\bar{y}^{k}].\end{array}\right.$ (2.8)
Writing (2.7) component-wisely, we get the relaxed and preconditioned Douglas-Rachford iteration as follows
$\left\{\begin{array}[]{ll}x^{k+1}=(1-\rho_{k})x^{k}+\rho_{k}x_{t}^{k+1},\\ y^{k+1}=(1-\rho_{k})y^{k}+\rho_{k}y_{t}^{k+1},\\ \bar{x}^{k+1}=\bar{x}^{k}+\rho_{k,x}[x_{\operatorname{test}}^{k+1}-x_{t}^{k+1}],\\ \bar{y}^{k+1}=\bar{y}^{k}+\rho_{k,y}[y_{\operatorname{test}}^{k+1}-y_{t}^{k+1}].\\ \end{array}\right.$ (RPDR)
Now we define the iteration operator RPDR as $\mathcal{T}_{R}$ such that $u^{k+1}:=\mathcal{T}_{R}u^{k}$. Note that PDR is a special case of RPDR when $I_{\varrho_{k}}=I$. Let us define
$\mathbf{M}_{\varrho_{k}}:=\mathbf{M}I^{-1}_{\varrho_{k}}=\operatorname{Diag}\left[\frac{1}{\sigma\rho_{k}}(N_{1}-I),\frac{1}{\tau\rho_{k}}(N_{2}-I),\frac{1}{\sigma\rho_{k,x}}I,\frac{1}{\tau\rho_{k,y}}I\right].$
We will present the following proposition, with the proof provided in the Appendix.
Proposition 2.
If ${u={(x,y,\bar{x},\bar{y})}}$ with $(x,y)\in\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$ and $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$, $\bar{y}=y-\tau\mathcal{K}x$, then we have
$\mathcal{L}(x_{\operatorname{test}}^{k+1},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k+1})\leq\frac{1}{2}(\|u^{k}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}}^{2}).$ (2.9)
@@
The iteration RPDR for problem 1.2 possesses the following properties:
@@
    A point $u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})$ is a fixed point of RPDR if and only if $(x^{*},y^{*})$ is a saddle point for (1.2) and $x^{*}+\sigma\mathcal{K}^{*}y^{*}=\bar{x}^{*},y^{*}-\tau\mathcal{K}x^{*}=\bar{y}^{*}$.
@@
Remark 1.
If we choose $I_{\varrho_{k}}:=\textrm{Diag}[I,I,\rho_{k,x}I,\rho_{k,y}I]$ in relaxed preconditioned Douglas-Rachford method (2.7), we get partial relaxed preconditioned Douglas-Rachford method (p-RPDR). Further, if we choose $N_{1}=I$, $N_{2}=I$, and $\rho_{k,x}=\rho_{k,y}$, we derive the original relaxed Douglas-Rachford splitting that was studied in [19], of which the weak convergence can be seen in [19, Theorem 7] with an ergodic convergence rate.
[…]
To clarify, we refer to the general RPDR as the fully relaxed preconditioned Douglas-Rachford method (f-RPDR) when $I_{\varrho_{k}}:=\textrm{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,x}I,\rho_{k,y}I]$, where $\rho_{k}\neq 1$, $\rho_{k,x}\neq 1$, and $\rho_{k,y}\neq 1$.
@@
$\mathcal{F}(x)=\langle\frac{1}{2}Qx-f,x\rangle,$ (2.10)
where $Q$ is a self-adjoint, linear, and bounded positive semi-definite operator and $f\in X$, we can get a more compact preconditioned Douglas–Rachford splitting method for quadratic problems, which is called PDRQ [10, 9, 7]. Denoting by $u=(x,y,\bar{y})\in\mathbb{U}_{Q}:=\mathbb{X}\times\mathbb{Y}\times\mathbb{Y}$, $N_{1}\succeq 0$, $N_{2}\succeq I$, the PDRQ iteration reads as,
@@
Without the updates for $\bar{x}^{k+1}$, the PDRQ iterations are more compact and efficient compared to PDR [10]. Moreover, if we choose $N_{2}=I$ and $N_{1}$ such that
@@
where $M_{Q}:X\rightarrow X$ is linear, continuous, self-adjoint and feasible meaning $M_{Q}\succeq T_{Q}:=\sigma Q+\sigma\tau\mathcal{K}^{*}\mathcal{K}$, then we can update $x^{k+1}$ first with the following classical preconditioned iteration
@@
The linear update of $y^{k+1}=\bar{y}^{k}+\tau\mathcal{K}x^{k+1}$ is then according to the second line in PDRQ [10, 9, 7]. Denoting
@@
Let $\mathcal{T}_{Q}$ denote the iteration operator associated with PDRQ, we can rewrite the iteration PDRQ as
@@
Now, we introduce the relaxed PDRQ method,
$u^{k+1}=(I-I_{q,\varrho_{k}})u^{k}+I_{q,\varrho_{k}}\mathcal{T}_{Q}u^{k},$ (2.15)
where $I_{q,\varrho_{k}}:=\textrm{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,y}I]$. Suppose $\rho_{k}$, $\rho_{k,y}$ are non-decreasing and have lower and upper limits $\rho_{l},\rho_{u}$, and $20<\rho_{l}<\rho_{u}<$ 2. We also set $u^{k+1}_{t}=\mathcal{T}_{{Q}}u^{k}:=(x^{k+1}_{t},y^{k+1}_{t},\bar{y}^{k+1}_{t})$ and
$y_{\operatorname{test}}^{k+1}=(I+\tau\partial\mathcal{G})^{-1}[2y_{t}^{k+1}-y^{k}],$
Writing (2.15) component-wisely, we get the relaxed and preconditioned Douglas-Rachford iteration for the quadratic problem as follows
$\left\{\begin{array}[]{ll}x^{k+1}=(1-\rho_{k})x^{k}+\rho_{k}x_{t}^{k+1},\\ y^{k+1}=(1-\rho_{k})y^{k}+\rho_{k}y_{t}^{k+1},\\ \bar{y}^{k+1}=\bar{y}^{k}+\rho_{k,y}[y_{\operatorname{test}}^{k+1}-y_{t}^{k+1}].\\ \end{array}\right.$ (RPDRQ)
We now define the iteration operator RPDRQ as $\mathcal{T}_{RQ}$ such that $u^{k+1}:=\mathcal{T}_{RQ}u^{k}$. Let $\mathbf{M}_{Q\varrho_{k}}:=\mathbf{M}_{Q}I^{-1}_{q,\varrho_{k}}=\operatorname{Diag}[\frac{1}{\sigma\rho_{k}}N_{1},\frac{1}{\tau\rho_{k}}(N_{2}-I),\tfrac{1}{\tau\rho_{k,y}}I]$. Similarly, we have the following proposition and lemma for the iteration sequence $\{u^{k}\}_{k}$ produced by RPDRQ.
@@
$\mathcal{L}(x_{\operatorname{test}}^{k+1},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k+1})\leq\frac{1}{2}(\|u^{k}-u\|_{\mathbf{M}_{Q\varrho_{k}}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}_{Q\varrho_{k}}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}_{Q}(2I-I_{q,\varrho_{k}})I^{-2}_{q,\varrho_{k}}}^{2}).$ (2.16)
@@
The iteration RPDRQ for 1.2 with $\mathcal{F}$ as in (2.10) possesses the following properties:
@@
Now, we introduce the restricted primal-dual gap function and the Bregman distance for the gap estimates. Suppose that the bounded set $\mathbb{B}:=\mathbb{B}_{1}\times\mathbb{B}_{2}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$ contains a saddle point. We give the following restricted primal-dual gap $\mathfrak{G}_{\mathbb{B}_{1}\times\mathbb{B}_{2}}$ associated with $\mathbb{B}$: